OpenStudy (anonymous):

Find the graph of the inequality y + 2

OpenStudy (anonymous):

ok so what do you do first

OpenStudy (anonymous):

the points are also (0,-3) and (4,-2)

OpenStudy (anonymous):

Find the slope?

OpenStudy (anonymous):

even though its 1/4

OpenStudy (anonymous):

how about distribute

OpenStudy (sparklestaraa):

Theres this amazing site called mathway, just type it in on google :D

OpenStudy (sparklestaraa):

from basic math to trigonometry, calculus and statistics, algebra, and graphing :D ,infinite and linear math :D

OpenStudy (anonymous):

i know i have used it before.

OpenStudy (sparklestaraa):

oh ok then try wolfram alpha

OpenStudy (sparklestaraa):

i would say put that into desmos graphing calculator :D

OpenStudy (sparklestaraa):

explain what </ means? is that greater than and equal to?

OpenStudy (anonymous):

its the >< with a line under it and its the equal to sign.

OpenStudy (anonymous):

you need to get it into y=mx+b format first

OpenStudy (sparklestaraa):

ohhhh ok :D yes than greater than and equal to

OpenStudy (sparklestaraa):

y + 2 </ 1/4 (x-4). rewrite this y plus 2 is greater than and equal to 1/4 multiplied by x minus 4

OpenStudy (sparklestaraa):

get y by itself y= </ 1/4 +2 (x-4) y= 2 and 1/4 (x-4)

OpenStudy (sparklestaraa):

good so far?

OpenStudy (sparklestaraa):

here's your graph because I'm learning the same things and really I'm not trying to make you know it on your own, i understand the struggle so here :P

OpenStudy (anonymous):

Thanks So are you a freshman/sophomore? @Sparklestaraa

OpenStudy (sparklestaraa):

freshman :D online school best thing ever :D

OpenStudy (anonymous):

Its better for some but not for me.

OpenStudy (sparklestaraa):

Yes yes true whats your struggles?

OpenStudy (anonymous):

ok so will it be dashed if its > or < and will it be a solid line if its \[\ge \le \]?

OpenStudy (anonymous):

right

OpenStudy (anonymous):

does it matter which way the > < sign is?

OpenStudy (anonymous):

@amanda7414

OpenStudy (anonymous):

@Tallan

OpenStudy (anonymous):

HI xD

OpenStudy (anonymous):

oops computer messed up didnt know you replied and it does i thoink